# CORTEX-AVD: CORner Case Testing & EXploration for Autonomous Vehicles Development

### v1

@@ Abstract
Autonomous Vehicles (AVs) aim to improve traffic safety and efficiency by reducing human error. However, ensuring AVs reliability and safety is a challenging task when rare, high-risk traffic scenarios are considered. These ’Corner Cases’ (CC) scenarios, such as unexpected vehicle maneuvers or sudden pedestrian crossings, must be safely and reliable dealt by AVs during their operations. But they are hard to be efficiently generated. Traditional CC generation relies on costly and risky real-world data acquisition, limiting scalability, and slowing research and development progress. Simulation-based techniques also face challenges, as modeling diverse scenarios and capturing all possible CCs is complex and time-consuming. To address these limitations in CC generation, this research introduces CORTEX-AVD, CORner Case Testing & EXploration for Autonomous Vehicles Development, an open-source framework that integrates the CARLA Simulator and Scenic to automatically generate CC from textual descriptions, increasing the diversity and automation of scenario modeling. Genetic Algorithms (GA) are used to optimize the scenario parameters in six case study scenarios, increasing the occurrence of high-risk events. Unlike previous methods, CORTEX-AVD incorporates a multi-factor fitness function that considers variables such as distance, time, speed, and collision likelihood. Additionally, the study provides a benchmark for comparing GA-based CC generation methods, contributing to a more standardized evaluation of synthetic data generation and scenario assessment. Experimental results demonstrate that the CORTEX-AVD framework significantly increases CC incidence while reducing the proportion of wasted simulations.
@@ I Introduction
AVs are safety-critical systems, as failures can have serious consequences, including environmental damage, and financial or life losses [5, 8, 9]. To mitigate AV safety risks, widely adopted industry standards, such as SAE J3016, which defines driving automation levels [10], and ISO 26262, which outlines functional safety requirements for road vehicles [11], provide essential guidelines for the development of AV [12]. These standards emphasize the need for rigorous validation processes to ensure AVs can handle a wide range of scenarios, including rare and unpredictable situations. Consequently, identifying and incorporating these challenging scenarios, often referred to as Corner Cases (CCs), is necessary for developing robust control algorithms capable of enhancing AV safety [13, 14].
CCs represent atypical scenarios that rarely occur in everyday driving, but can lead to severe consequences if not handled properly. Examples include unexpected vehicle behavior, sudden pedestrian crossings, environmental factors that perturb sensors, or obstacles on the road [15]. However, some current approaches are heavily based on real-world data collection, which is costly, time-consuming, and inherently limited in capturing the diversity of rare events [16, 5, 17, 18]. Physical tests, such as those conducted at facilities such as Mcity University in Michigan [19] or Waymo’s Castle [20, 21], offer controlled environments but fail to cover the full spectrum of potential CCs [15]. To mitigate these issues, researchers are refining techniques to identify and generate CC, while using simulations to improve the safety of AV [22].
On the other hand, simulated environments offer a promising alternative for controlled and repeatable AV testing in high-risk scenarios [23, 24, 25, 26, 27, 28], yet generating diverse CCs remains a complex task [15]. Current methodologies often rely on labor-intensive scenario modeling, expert domain knowledge, or proprietary tools, limiting accessibility and slowing progress in developing robust control algorithms [29, 30, 31, 32]. Furthermore, the black-box nature of Deep Learning (DL), the predominant model in AV decision-making [5, 22, 1], complicates scenario validation [15, 12], making it difficult to ensure safe behavior under CC conditions [33, 34, 22], which raises concerns about transparency and public trust in AV technology [35, 36]. While synthetic data generation and simulated environments have driven safety improvements, academic research remains fragmented, delayed by the absence of standardized benchmarks or unified testing frameworks, which limits meaningful collaboration between academia and industry [29, 30, 31, 32]. Given the safety-critical nature of AVs, establishing open and standardized practices for CC generation is necessary to advance AV safety [30].
@@
  • Large-scale case study demonstrating improved risk scenario refinement and simulation validity.
This study is structured into 6 sections. Section II presents the related work. Then, Section III presents the methodology. Section IV presents the results. Finally, Section V provides a discussion and Section VI the concluding remarks on the findings.
@@ II Related Work
Although recent studies highlight the effectiveness of GA in CC generation, most experiments are limited to a few thousand simulations and rely on single-objective functions based on narrow metrics such as time or distance [62, 65, 63, 66, 56, 57, 68, 58, 59, 60, 61]. These approaches may neglect variations of CCs, as they may fail to adequately capture the complexity of real-world driving scenarios. By contrast, in other fields, multi-objective optimization enables a more comprehensive search, offering a broader exploration of potential solutions [69, 70]. Only few studies using GAs rely on multi-objective optimization techniques [65, 66, 67, 68], allowing for the balancing of factors like safety, efficiency, and complexity.
@@
Beyond the limitations previously discussed, these methods share common challenges. Techniques like [38, 54, 57, 64] face high computational demands, requiring specialized hardware and limiting scalability in complex scenarios. Some of these approaches struggle with simulation determinism and the realism of agent behavior [67, 59]. Thus, to address these challenges and enhance the applicability of these methods, further research is needed to improve computational efficiency, scalability, and realism, with multi-factor approaches and standardized evaluation criteria to enable fair comparisons across studies and assess performance consistently.
@@ III The CORTEX-AVD Framework
This section presents the proposed framework for automatically generating CCs to support the development of robust vehicle control systems based on high-risk driving scenarios. The following subsections describe the simulation infrastructure (III-A), the framework used to implement GA (III-B), the experimental design employed in the case study (III-C), and the metrics used to evaluate and compare experiments (III-D).
@@ III-A Simulation Infrastructure
The Scenic is a probabilistic programming language to define complex traffic scenarios by specifying spatial and temporal relationships between agents and physical entities, with constraints that can range from strict to more relaxed conditions [72]. Its concise syntax simplifies the generation of diverse and realistic driving scenarios [29]. In this study, Scenic was chosen for its efficiency in generating scenarios from formal descriptions, which are then integrated into Carla simulations to create diverse testing environments that meet specified feasibility criteria [72].
@@ III-B Parameters Refinement
In this study, a Genetic Algorithm (GA) is employed to optimize the search for parameter vectors $\vec{x}\in\mathbb{R}^{n}$ within Scenic scripts, aiming to generate high-risk scenarios such as accidents and near-misses. Inspired by natural selection, GAs have proven effective in navigating complex, nonlinear parameter spaces to identify critical conditions in dynamic systems [69, 70]. Traditional methods, such as manually mapping or exhaustively listing potential accident scenarios, are impractical due to the system’s high dimensionality and nonlinear behavior. In contrast, GAs are particularly well-suited for this task due to their ability to explore large search spaces without relying on gradient information. Unlike gradient-based optimization methods, which require costly derivative computations and struggle with non-differentiable or noisy objective functions, GAs without relying on gradients, significantly reducing computational costs while maintaining robustness [73]. Their scalability and global search capabilities enable efficient convergence to high-risk configurations, even in large-scale simulations. Moreover, GAs adapt seamlessly as new elements are introduced during scenario evolution. As shown on the right side of Figure 1, this optimization process forms a feedback loop where the GA adjusts the input parameters $\vec{x}$, resulting in updated simulation outcomes $f(\vec{x})=\{y_{1},y_{2},\dots,y_{k}\}$. This loop illustrates how the framework dynamically refines scenario configurations to uncover edge cases and critical testing conditions.
@@ III-B2 Fitness Function
By combining insights from the literature with the data collected from the vehicles during the simulations, hypotheses were formed regarding metrics useful for the objective function (Table I) [5, 74, 53, 75, 76, 77]. Based on these references, the multi-factor objective function was constructed using variables frequently associated with driving risk: Collision (C), Minimum Distance between vehicles (MD), Distance at Maximum Approach Speed (D_MS), and Time-to-Collision (TTC) at the moment of Maximum Approach Speed (TTC_MS) - as shown in Table II. The function returns a value in the range $[0,22]$, where higher scores indicate greater scenario risk. The function was evaluated through experiments to verify its ability to distinguish scenarios with different risk levels, with results presented in Section IV indicating that it contributed meaningfully to the search process.
@@
Additionally, due to the Scenic’s sampling nature, certain parameter sets may produce invalid or non-executable test cases. These are assigned a risk score of -1 and excluded from the GA process. During scenario generation, only valid cases are retained for selection, crossover, and mutation, ensuring that the optimization operates solely on executable simulations.
@@ III-B3 GA Parameter Tuning
The GA uses three operations, selection, crossover, and mutation, with their probabilities defined as $\mu_{s}$, $\mu_{c}$, and $\mu_{m}$, respectively, such that $\mu_{s}+\mu_{c}+\mu_{m}=1$. During each generation, while the new population has not yet reached the specified population size, an operation is chosen by drawing a value randomly from a uniform distribution in the range $[0,1]$ which is then mapped to one of the three operations based on their probabilities. If the value falls within $[0,\mu_{s}]$, elitism is applied, selecting the best individual not already chosen from the prior generation, ensuring no repetition [78]. If the value is in $[\mu_{s},(\mu_{s}+\mu_{c})[$, single-point crossover occurs between two parents, producing two new individuals [78]. Finally, if the value is in $[(\mu_{s}+\mu_{c}),1]$, random mutation is applied within a predefined range [78]. For the experiments, $\mu_{s}=0.1$, $\mu_{c}=0.8$, and $\mu_{m}=0.1$ were chosen based on empirical results. A 10% mutation rate was adopted to increase variability in the search space while preserving high-quality solutions from previous generations [79].
III-C Experimental Design
The case study aimed to validate the hypothesis H1: simulations generated by GA have a higher likelihood of collision or near-collision occurrences compared to random sampling method. Intersection scenarios involving two moving vehicles were selected based on the NHTSA’s 2011–2015 light vehicle pre-crash statistics [80], using scenarios with high accident incidence [81]. These scenarios, listed in Table III and illustrated in Figure 2, were chosen to generate relevant data on high-risk events, supported by the module that generates tuned parameters.
Figure 2: Set of test scenarios used in the case study according to Table III
Figure 3: Case study infrastructure
@@ III-C Experimental Design
A 2 x 2 Vehicle A performs a crossing
Vehicle B (same lane, opposite direction) performs a left turn
B 2 x 2 Vehicle A performs a left turn
Vehicle B (perpendicular lane) performs a crossing
C 2 x 2 Vehicle A performs a crossing or left turn
Vehicle B (perpendicular lane) performs a left turn
D 2 x 2 Vehicle A performs a right turn
Vehicle B (same lane, opposite direction) performs a left turn
E 2 x 2 Vehicle A performs a right turn
Vehicle B (perpendicular lane) crosses in the same direction as A
F 3 Vehicle A performs a left turn
Vehicle B (perpendicular lane) crosses in the same direction as A
@@
To ensure a standardized comparison between the GA and random approaches, the same number of simulations was used for each. Specifically, for each generation, 100 scenarios were generated by the GA and 100 scenarios were produced by sampling random parameters, mirroring the same simulation structure and volume. This setup ensured that both methods operated under equivalent conditions for comparative analysis.
III-D Evalutation Metrics
IV Results
@@ IV-B Number of Collisions
IV-C Minimum Distances
The MD metric evaluates the closest proximity between vehicles to assess risk during simulations, with the level of risk depending on the specific actions and context of the vehicles involved. In normal traffic, vehicles may approach each other without posing significant risk, but in these experiments, which forces confrontations in conflict zones during crossing and turning actions, the MD reflects the aggressiveness of their approach [83]. Figures 8 and 9 show the average minimum approach distances, with the first considering all valid scenarios (MDG) and the second excluding simulations that resulted in collisions (MDEC) to focus on near-accident situations. As seen with previous metrics, GA scenarios exhibit a pattern of increasing risk, with MDs decreasing over generations, while Random approach averages remain near-uniform across generations. The trends are well represented by Figure 4 and the third and fourth rows of Figure 5, highlighting the superiority of GA in all scenarios, despite the regression of MD in scenario D in the final generations. Notably, in the last 10 generations [21-30], MDG showed GA’s superiority across all scenarios: A (GA: 69.4 — R: 138.9), B (GA: 121.5 — R: 187.5), C (GA: 101.4 — R: 162.5), D (GA: 144.6 — R: 180.8), E (GA: 143.9 — R: 213.6), and F (GA: 166.8 — R: 200.7). In MDEC, even with collisions removed, showed similar results, though the average distances were slightly higher for all scenarios: A (GA: 77.9 — R: 161.2), B (GA: 141.8 — R: 217.6), C (GA: 122.6 — R: 178.2), D (GA: 160.9 — R: 195.7), E (GA: 154.7 — R: 224.6), and F (GA: 191.5 — R: 236.1). Thus, GA consistently generated riskier scenarios with decreasing MDs, even in MDEC, reinforcing its ability to explore near-accident conditions more effectively than Random method.

### v3

# CORTEX-AVD: A Framework for CORner Case Testing and EXploration in Autonomous Vehicle Development

@@ Abstract
Autonomous Vehicles (AVs) aim to improve traffic safety and efficiency by reducing human error. However, ensuring AVs reliability and safety is a challenging task when rare, high-risk traffic scenarios are considered. These ’Corner Cases’ (CC) scenarios, such as unexpected vehicle maneuvers or sudden pedestrian crossings, must be safely and reliably dealt by AVs during their operations. But they are hard to be efficiently generated. Traditional CC generation relies on costly and risky real-world data acquisition, limiting scalability, and slowing research and development progress. Simulation-based techniques also face challenges, as modeling diverse scenarios and capturing all possible CCs is complex and time-consuming. To address these limitations in CC generation, this research introduces CORTEX-AVD, an open-source framework that integrates the CARLA Simulator and Scenic to automatically generate CC from textual descriptions, increasing the diversity and automation of scenario modeling. Genetic Algorithms (GA) are used to optimize the scenario parameters in six case study scenarios, increasing the occurrence of high-risk events. Unlike previous methods, CORTEX-AVD incorporates a multi-factor fitness function that considers variables such as distance, time, speed, and collision occurrence. Additionally, the study provides a benchmark for comparing GA-based CC generation methods, contributing to a more standardized evaluation of synthetic data generation and scenario assessment. Experimental results demonstrate that the CORTEX-AVD framework significantly increases CC incidence while reducing the proportion of wasted simulations.
@@ I Introduction
AVs are safety-critical systems, as failures can have serious consequences, including environmental damage, and financial or life losses [5, 8, 9]. To mitigate AV safety risks, widely adopted industry standards, such as SAE J3016, which defines driving automation levels [10], and ISO 26262, which outlines functional safety requirements for road vehicles [11], provide essential guidelines for the development of AV [12]. These standards emphasize the need for rigorous validation processes to ensure AVs can handle a wide range of scenarios, including rare and unpredictable situations. Consequently, identifying and incorporating these challenging scenarios, often referred to as Corner Cases (CCs), is necessary for developing robust control algorithms capable of enhancing AV safety [13, 14, 15].
CCs represent atypical scenarios that rarely occur in everyday driving, but can lead to severe consequences if not handled properly. Examples include unexpected vehicle behavior, sudden pedestrian crossings, environmental factors that perturb sensors, or obstacles on the road [16]. However, some current approaches are heavily based on real-world data collection, which is costly, time-consuming, and inherently limited in capturing the diversity of rare events [17, 5, 18, 19]. Physical tests, such as those performed at testing facilities like Mcity, located at the University of Michigan [20], or Waymo’s Castle [21, 22], offer controlled environments but fail to cover the full spectrum of potential CCs [16]. To mitigate these issues, researchers are focusing on simulation-based methods to identify and generate CCs more efficiently [23, 15].
Simulated environments offer a promising alternative for controlled and repeatable AV testing in high-risk scenarios [24, 25, 26, 27, 28, 29], yet generating diverse CCs remains a complex task [16]. Current methodologies often rely on labor-intensive scenario modeling, expert domain knowledge, or proprietary tools, limiting accessibility and slowing progress in developing robust control algorithms [30, 31, 32, 33]. Furthermore, the black-box nature of Deep Learning (DL), the predominant model in AV decision-making [5, 23, 1], complicates scenario validation [16, 12], making it difficult to ensure safe behavior under CC conditions [34, 35, 23]. These issues raise concerns about transparency and public trust in AV technology [36, 37]. Although synthetic data generation and simulated environments have driven safety improvements, academic research remains fragmented, delayed by the absence of standardized benchmarks or unified testing frameworks, which limits meaningful collaboration between academia and industry [30, 31, 32, 33]. Given the safety-critical nature of AVs, establishing open and standardized practices for CC generation is necessary to advance AV safety [31, 15].
@@
  • A case study demonstrating improved risk scenario refinement and simulation validity.
This study is structured into 7 sections. Section II presents the related work. Then, Section III presents the methodology. Section IV presents the proposed case study. Section V presents the results. Finally, Section VI provides a discussion and Section VII the concluding remarks on the findings.
@@ II Related Work
Although recent studies highlight the effectiveness of GA in CC generation, most experiments are limited to a few thousand simulations and rely on single-objective functions based on narrow metrics such as time or distance [63, 66, 64, 67, 57, 58, 69, 59, 60, 61, 62]. These approaches may neglect variations of CCs, as they may fail to adequately capture the complexity of real-world driving scenarios. By contrast, in other fields, multi-objective optimization enables a more comprehensive search, offering a broader exploration of potential solutions [70, 71]. However, in the context of AV CC generation, only a few studies using GAs rely on multi-objective optimization techniques [66, 67, 68, 69], allowing for the balancing of factors like safety, efficiency, and complexity.
@@
Beyond the limitations previously discussed, these methods share common challenges. Techniques like [39, 55, 58, 65] face high computational demands, requiring specialized hardware and limiting scalability in complex scenarios. Some of these approaches struggle with simulation determinism and the realism of agent behavior [68, 60]. Thus, to address these challenges and enhance the applicability of these techniques, further research is needed to improve computational efficiency, scalability, and realism, with multi-factor approaches and standardized evaluation criteria to enable fair comparisons across studies and assess performance consistently.
@@ III The CORTEX-AVD Framework
This section presents the proposed framework for automatically generating CCs to support the development of robust vehicle control systems based on high-risk driving scenarios. The following subsections describe the simulation infrastructure (III-A), the framework used to implement GA (III-B), and the metrics used to evaluate and compare experiments (III-C).
@@ III-A Simulation Infrastructure
The Scenic is a probabilistic programming language to define complex traffic scenarios by specifying spatial and temporal relationships between agents and physical entities, with constraints that can range from strict to more relaxed conditions [73]. Its concise syntax simplifies the generation of diverse and realistic driving scenarios [30]. Specifically, Scenic was selected for its efficiency in generating scenarios from formal descriptions within Carla simulator [73]. However, the framework is not limited to Scenic and can be adapted to support other high-level, textual scenario description languages.
@@ III-B Parameters Refinement
In the framework, a Genetic Algorithm (GA) is employed to optimize the search for parameter vectors $\vec{x}\in\mathbb{R}^{n}$ within Scenic scripts, aiming to generate high-risk scenarios such as accidents and near-misses. Inspired by natural selection, GAs have proven effective in navigating complex, nonlinear parameter spaces to identify critical conditions in dynamic systems [70, 71]. Traditional methods, such as manually mapping or exhaustively listing potential accident scenarios, are impractical due to the system’s high dimensionality and nonlinear behavior. In contrast, GAs are particularly well-suited for exploring large search spaces without relying on gradient information, avoiding costly derivative computations and handling non-differentiable or noisy objective functions with greater robustness [74]. Their scalability and global search capabilities enable efficient convergence to high-risk configurations, even in large-scale simulations. Moreover, GAs adapt seamlessly as new elements are introduced during scenario evolution. As shown on the right side of Figure 1, this optimization process forms a feedback loop where the GA adjusts the input parameters $\vec{x}$, resulting in updated simulation outcomes $f(\vec{x^{\prime}})=\{y_{1}^{\prime},y_{2}^{\prime},\dots,y_{k}^{\prime}\}$. This loop illustrates how the framework dynamically refines scenario configurations to uncover edge cases and critical testing conditions.
@@ III-B2 Fitness Function
By combining insights from the literature with the data collected from the vehicles during the simulations, hypotheses were formed regarding metrics useful for the objective function (Table I) [5, 75, 54, 76, 77, 78]. Based on these references, the multi-factor objective function was constructed as the sum of variables frequently associated with driving risk: Collision (C), Minimum Distance between vehicles (MD), Distance at Maximum Approach Speed (D_MS), and Time-to-Collision (TTC) at the moment of Maximum Approach Speed (TTC_MS) (Table II). The function returns a value in the range $[0,22]$, where higher scores indicate greater scenario risk. The function was evaluated through experiments to verify its ability to distinguish scenarios with different risk levels, with results presented in Section V indicating that it contributed meaningfully to the search process.
@@
Additionally, due to the Scenic’s sampling nature, some parameter sets may produce invalid or non-executable test cases. These are assigned a risk score of -1 and excluded from the GA process. During scenario generation, only valid cases are retained for selection, crossover, and mutation, ensuring that the optimization operates solely on executable simulations.
@@ III-B3 GA Parameter Tuning
The GA employs three operations - selection, crossover, and mutation - each associated with a probability: selection rate ($\mu_{s}$), crossover rate ($\mu_{c}$), and mutation rate ($\mu_{m}$), respectively, such that $\mu_{s}+\mu_{c}+\mu_{m}=1$. During each generation, while the new population has not yet reached the target size, an operation is selected by randomly sampling a value from a uniform distribution in the range $[0,1]$ and mapping it to one of the three operations based on their probabilities. If the value falls within $[0,\mu_{s}]$, elitism is applied, selecting the best individual not already chosen from the prior generation, ensuring no repetition [79]. If it lies in $[\mu_{s},(\mu_{s}+\mu_{c})[$, single-point crossover occurs between two parents, producing two new individuals [79]. Finally, if the value is in $[(\mu_{s}+\mu_{c}),1]$, random mutation is applied within a predefined range [79]. For the experiments, $\mu_{s}=0.1$, $\mu_{c}=0.8$, and $\mu_{m}=0.1$ were selected based on empirical results. A 10% mutation rate was adopted to increase variability in the search space while preserving high-quality solutions from earlier generations [80].
III-C Evalutation Metrics
The combination of simulation infrastructure, scenario optimization through GA, and structured evaluation metrics constitutes the foundation of the CORTEX-AVD framework. To assess the applicability and performance of this approach, the next section presents a case study involving its deployment in a controlled experimental setting.
IV Case Study
The case study aimed to validate the hypothesis H1: simulations generated by GA have a higher likelihood of collision or near-collision occurrences compared to random sampling method. Intersection scenarios involving two moving vehicles were selected based on the NHTSA’s 2011–2015 light vehicle pre-crash statistics [81], using scenarios with high accident incidence [82]. These scenarios, listed in Table III and illustrated in Figure 2, were used to generate relevant data on high-risk events, supported by the module responsible for tuning parameters.
Figure 2: Illustration of the NHTSA intersection scenarios used in the case study, corresponding to those listed in Table III
@@ IV Case Study
A 2 x 2 Vehicle 1 performs a crossing
Vehicle 2 (same lane, opposite direction) performs a left turn
B 2 x 2 Vehicle 1 performs a left turn
Vehicle 2 (perpendicular lane) performs a crossing
C 2 x 2 Vehicle 1 performs a crossing or left turn
Vehicle 2 (perpendicular lane) performs a left turn
D 2 x 2 Vehicle 1 performs a right turn
Vehicle 2 (same lane, opposite direction) performs a left turn
E 2 x 2 Vehicle 1 performs a right turn
Vehicle 2 (perpendicular lane) crosses in the same direction as A
F 3 Vehicle 1 performs a left turn
Vehicle 2 (perpendicular lane) crosses in the same direction as A
@@
To ensure a standardized comparison between the GA and random approaches, the same number of simulations was used for each. Specifically, for each generation, 100 scenarios were generated by the GA and 100 scenarios were produced by sampling random parameters, ensuring symmetry in simulation structure and volume. This setup ensured that both methods operated under equivalent conditions for comparative analysis.
Figure 3: Case study infrastructure
V Results
@@ V-B Number of Collisions
V-C Minimum Distances
The MD metric evaluates the closest proximity between vehicles to assess risk during simulations, with the level of risk depending on the specific actions and context of the vehicles involved. In normal traffic, vehicles may approach each other without posing significant risk, but in these experiments, which forces confrontations in conflict zones during crossing and turning actions, the MD reflects the aggressiveness of their approach [84]. Figures 8 and 9 show the average minimum approach distances (in centimeters), with the first considering all valid scenarios (MDG) and the second excluding simulations that resulted in collisions (MDEC) to focus on near-accident situations. As seen with previous metrics, GA scenarios exhibit a pattern of increasing risk, with MDs decreasing over generations, while Random approach averages remain near-uniform across generations. The trends are well represented by Figure 4 and the third and fourth rows of Figure 5, highlighting the superiority of GA in all scenarios, despite the regression of MD in scenario D in the final generations. Notably, in the last 10 generations [21-30], MDG showed GA’s superiority across all scenarios: A (GA: 69.4 — R: 138.9), B (GA: 121.5 — R: 187.5), C (GA: 101.4 — R: 162.5), D (GA: 144.6 — R: 180.8), E (GA: 143.9 — R: 213.6), and F (GA: 166.8 — R: 200.7). In MDEC, even with collisions removed, showed similar results, though the average distances were slightly higher for all scenarios: A (GA: 77.9 — R: 161.2), B (GA: 141.8 — R: 217.6), C (GA: 122.6 — R: 178.2), D (GA: 160.9 — R: 195.7), E (GA: 154.7 — R: 224.6), and F (GA: 191.5 — R: 236.1). Thus, GA consistently generated riskier scenarios with decreasing MDs, even in MDEC, reinforcing its ability to explore near-accident conditions more effectively than Random method.